# MMG: Mutual Information Estimation via the MMSE Gap in Diffusion

### v1

@@ Abstract
Mutual information (MI) is one of the most general ways to measure relationships between random variables, but estimating this quantity for complex systems is challenging. Denoising diffusion models have recently set a new bar for density estimation, so it is natural to consider whether these methods could also be used to improve MI estimation. Using the recently introduced information-theoretic formulation of denoising diffusion models, we show the diffusion models can be used in a straightforward way to estimate MI. In particular, the MI corresponds to half the gap in the Minimum Mean Square Error (MMSE) between conditional and unconditional diffusion, integrated over all Signal-to-Noise-Ratios (SNRs) in the noising process. Our approach not only passes self-consistency tests but also outperforms traditional and score-based diffusion MI estimators. Furthermore, our method leverages adaptive importance sampling to achieve scalable MI estimation, while maintaining strong performance even when the MI is high.
@@ 1 Introduction
Estimating Mutual Information (MI) from samples is a fundamental problem with widespread applications. Methods based on local density estimation (Kraskov et al., 2004; Pál et al., 2010; Gao et al., 2015) have been displaced by variational approaches using neural networks (Poole et al., 2019) to estimate lower bounds on MI (Belghazi et al., 2018b; Nguyen et al., 2010). Unfortunately, these approaches may have sample complexity or variance which scale exponentially with the true MI (Gao et al., 2015; McAllester & Stratos, 2020; van den Oord et al., 2018). In practical scenarios, the MI between two variables is usually unknown, making reliable estimation challenging. To tackle this, Song & Ermon (2019) introduced three self-consistency experiments designed to evaluate the robustness of MI estimators. For a more standardized and comprehensive evaluation, Czyż et al. (2023) developed a benchmark consisting of 40 synthetic datasets derived from diverse distributions, providing a consistent basis for assessing the performance of different MI estimators.
Denoising diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Kingma et al. (2021); Kong et al. (2022) have dramatically improved the modeling of complex distributions, igniting a new industry for AI art generation Rombach et al. (2022); Ramesh et al. (2022); Saharia et al. (2022). Recently, Franzese et al. (2024) introduced a MI estimator, MINDE, based on score-based diffusion models. This estimator not only achieved an 87.5% estimation success rate on the benchmark of Czyż et al. (2023) but also passed the self-consistency tests. However, its reliance on accurately approximating the log-density gradient can be a challenging intermediate step. This motivates a more direct formulation, connecting MI to the denoising objective itself rather than its gradient.
In this paper, we exploit the recently discovered connections between information theory and diffusion models to derive an elegant and effective approach to MI estimation (Guo et al., 2005; Kong et al., 2022). Our contributions are as follows:
  • We show that conditional density estimation and mutual information estimation can both be written exactly in terms of the global optimum of a denoising objective. Conditional density estimation corresponds to a gap between MMSEs (Kong et al., 2022) for conditional and unconditional denoising diffusion models, and mutual information is the expected gap over all data, leading to the Mutual Information Estimation via the MMSE Gap in Diffusion (MMG) estimator.
@@ 2 Background
Let $p(\bm{z}_{\gamma}|{\bm{x}})$ be a Gaussian noise channel with $\bm{z}_{\gamma}=\sqrt{\gamma/(1+\gamma)}{\bm{x}}+\sqrt{1/(1+\gamma)}\bm{\epsilon}$ and $\bm{\epsilon}\sim\mathcal{N}(0,\mathbb{I})$, where $\gamma$ represents the Signal-to-Noise Ratio (SNR). The unknown data distribution is $p({\bm{x}})$. The MMSE refers to the minimum mean square error for recovering $\bm{x}$} in this noisy channel (Kong et al., 2022).
@@
From Kong et al. (2022) we see that log likelihood can be written exactly in terms of an expression that depends only on the MMSE solution to the Gaussian denoising problem.
@@ 3.1 Mutual Information Estimation
Conditioning on $y$ can only decrease the MMSE Wu & Verdú (2011), so the mutual information is non-negative as expected. While Eq. 7 appears to be novel, a version of Eq. 8 appeared in Guo et al. (2005). This result holds for discrete or continuous data Guo et al. (2005), or even mixed continuous and discrete, a particularly challenging problem Gao et al. (2017). We propose to use recent advances in denoising diffusion modeling to approximate the right-hand side to achieve better mutual information estimates.
@@ 3.2 Model Training
We parametrize $\hat{\bm{x}}_{\theta}(\bm{z}_{\gamma},\gamma,y)$ as some neural network that is trained with a mean square error loss to recover the signal, $\bm{x}$}, from noise and the auxiliary signal, $y$, across different SNRs, $\gamma$ (see Figure 2). In principle we should train a separate neural network that recovers data from noise without conditioning on $y$. However, we can borrow from existing architectures which train a single network for both conditional and unconditional denoising Ho & Salimans (2021). Conditional diffusion models have shown compelling results in Nichol et al. (2021); Ramesh et al. (2022); Saharia et al. (2022). During training we simply replace $y$ with a null value, $y=\emptyset$, some fraction of the time so that the same (conditional) network can learn to denoise in the absence of conditioning.
In practice, we can’t guarantee that the neural network finds the true global optimum, the MMSEs in Eq. 8. Because MMSEs appear with both signs in that expression, we can’t even guarantee that our network gives an upper or lower bound. While this is unfortunate, conventional wisdom suggests that sufficiently expressive neural networks do converge to global optima (Du et al., 2019; Jacot et al., 2018).
@@ 4.1 Adaptive Importance Sampling
As in Kong et al. (2022), we estimate integrals using importance weighted Monte Carlo estimators, with log-SNR values chosen from a logistic distribution with some location, $\mu$ and scale, $\sigma$. Since the shape of the integration area ($\Delta$ area in Fig. 2) is data-dependent, a fixed sampling distribution can be inefficient. For our adaptive variants, we therefore optimize this distribution for each specific task.
To find the data-adapted parameters $(\mu,\sigma)$, we employ a two-stage procedure. First, we train a preliminary model and analyze its conditional MMSE curve, $\text{MMSE}_{x|y}$. Inspired by the use of error landmarks in density estimation (Kong et al., 2022), our heuristic identifies the critical transition region of this curve, where the denoiser becomes effective. Specifically, we define the parameters as follows (where $d$ is the data dimensionality):
@@ 4.2 The Orthogonal Principle
To further enhance estimator stability, we incorporate the orthogonality principle from Kong et al. (2023). The principle states that for optimal denoisers, the gap in MMSE is precisely the expected squared distance between the conditional and unconditional estimators. This provides an alternative way to express the MI integrand. Formally, let $\hat{{\bm{x}}}(\bm{z}_{\gamma})=\mathbb{E}[{\bm{x}}|\bm{z}_{\gamma}]$ and $\hat{{\bm{x}}}(\bm{z}_{\gamma},y)=\mathbb{E}[{\bm{x}}|\bm{z}_{\gamma},y]$. The identity at a fixed SNR is:
@@
After expanding the left-hand side, the cross terms vanish due to the orthogonality principle Kong et al. (2023), resulting in the expression on the right-hand side. This allows us to substitute the original integrand in our MI formula (Eq. 8) with the term on the right-hand side. The orthogonal MI estimator, in its practical importance-sampled form, is therefore:
@@ 5.1 MI Estimation Benchmark
We evaluate on the benchmark of (Czyż et al., 2023), which combines base distributions (Uniform, Normal with dense or sparse correlation, and long-tailed Student-t) with MI-preserving nonlinear transformations (Half-Cube, Asinh, Swiss-roll, Spiral). This setup introduces high dimensionality, heavy tails, sparsity, and non-linear geometry. We compare MMG against neural estimators, such as mine (Belghazi et al., 2018a), Infonce (Oord et al., 2018),nwj (Nguyen et al., 2007), doe (McAllester & Stratos, 2020) and MINDE Franzese et al. (2023).
@@
Table 1: Low MI estimation over 10 seeds using N = 10k test samples against ground truth (GT) (Czyż et al., 2023). All methods were trained with 100k samples. Color indicates relative negative (red) and positive bias (blue). Blank entries indicate that an estimator experienced numerical instabilities. List of abbreviations ( Mn: Multinormal, St: Student-t, Nm: Normal, Hc: Half-cube, Sp: Spiral)
@@ 5.2 High MI Benchmark
To test the robustness and limits of our estimator, we extend the high-MI study from MINDE (Franzese et al., 2024), pushing their experimental setup from a range of MI $\leq$ 5 into the significantly more challenging regime of MI $\in[10,15]$. Following their protocol, we use a sparse $3\times 3$ Multinormal setup and its two MI-preserving transforms (Half-cube, Spiral). Figure 3 reports the mean $\pm$ one standard deviation over 10 seeds.
On the original and Half-cube cases (Fig. 3a&b), MMG-adaptive remains the most accurate as MI grows, while other estimators falter due to distinct sources of error. MINDE struggles significantly as its score-matching objective requires approximating the sharp, high-frequency score functions of high-MI distributions, a task where neural networks’ spectral bias leads to significant underestimation Rahaman et al. (2019). In contrast, our orthogonal variants exhibit a different limitation: a systematic conservative bias. This occurs because the orthogonal estimator measures the distance between neural network approximations of the denoisers, and the distance between these "smoothed-out" approximations is inherently smaller than the true distance between the optimal denoisers. This underestimation bias becomes more pronounced as the true MI gap grows larger.
@@ 5.3 Consistency Test
We conduct self-consistency tests inspired by Song & Ermon (2019) to evaluate the properties of MMG using high-dimensional real-world data, specifically samples from the MNIST dataset (28×28 resolution) (Deng, 2012). Let $A$ represent an image and $B_{r}$ denote the image consisting of the top $r$ rows of $A$. These tests are designed to verify whether the estimators adhere to fundamental properties of MI through three subtasks: Baseline Test, Data-Processing Test, and Additivity Test for two independent images sampled from dataset. Ideally, as $r$ increases, $\frac{I(A;B_{r})}{I(A;B)}$ should monotonically approach 1, $\frac{I(A;[B_{r+k},B_{r}])}{I(A;B_{r+k})}$ should consistently equal 1, and $\frac{I([A^{1},A^{2}];[B^{1}_{r},B^{2}_{r}])}{I(A^{1};B^{1}_{r})}$ should consistently equal 2. To ensure a fair comparison with MINDE (Franzese et al., 2024), we aligned our experimental settings and parameters and tested MMG using five random seeds. The results of the three tests are presented in Figure 4. Overall, MMG performed well and successfully passed all tests.
@@ 6 Related Work
Estimating mutual information (MI) from samples is a central challenge in machine learning. Traditional non-parametric approaches, such as methods based on data binning or kernel density estimation (KDE), struggle with the curse of dimensionality. More advanced estimators based on k-nearest neighbors have shown significant improvements (Kraskov et al., 2004; Pál et al., 2010), but can still be challenged by complex, high-dimensional data (Gao et al., 2015). Consequently, the modern paradigm is dominated by neural network-based approaches that optimize a variational bound on the MI.
@@
Most recent neural MI estimators are built upon variational lower bounds derived from f-divergences. The pioneering work, MINE (Belghazi et al., 2018b), leverages the Donsker-Varadhan representation of the KL-divergence, spurring related estimators like the variance-reduced SMILE (Song & Ermon, 2019) and DoE (McAllester & Stratos, 2020). A particularly successful family is based on contrastive learning, where estimators like InfoNCE (van den Oord et al., 2018) train a critic to distinguish between joint and marginal samples (Poole et al., 2019). However, these methods face significant limitations: their sample complexity can scale exponentially with the true MI, often being capped by the logarithm of the batch size (McAllester & Stratos, 2020). This difficulty can be viewed as a "density chasm": the marginal distribution $p(x)p(y)$ is often a poor proposal for the joint $p(x,y)$, leading to high-variance estimates, especially in high-MI settings (Rhodes et al., 2020; Brekelmans et al., 2021).
@@
Denoising diffusion models offer a natural and powerful framework to bridge this density chasm. They define a continuous process that transforms a complex data distribution into a simple tractable one, providing a path of intermediate distributions. The potential of this framework for MI estimation was recently demonstrated by MINDE (Franzese et al., 2024), which connects MI to the difference between conditional and unconditional score functions ($\nabla_{{\bm{x}}}\log p({\bm{x}})$). Our work, MMG, builds on a different and more direct connection (Guo et al., 2005; Kong et al., 2022). Instead of relying on learned score functions, we show that MI corresponds exactly to the integrated gap between the Minimum Mean Square Error (MMSE) of conditional and unconditional denoising. This formulation connects MI directly to the denoising objective itself, rather than its gradient, providing an elegant and potentially more robust pathway for estimation.
@@ Appendix A Derivation of the Orthogonal Principle
This section provides a brief derivation for the orthogonal principle (Equation 11) used in our estimator, following the original work of Kong et al. (2023).
@@ Appendix C Implementation Details
We follow the implementation of Franzese et al. (2023) which uses stacked multi-layer perception (MLP) with skip connections. We adopt a simplified version of the same network architecture: this involves three Residual MLP blocks. We use the Adam optimizer (Kingma, 2014) for training and Exponential moving average (EMA) with a momentum parameter $m=0.999$. We use the ReduceLROnPlateau scheduler with a patience of 200 epochs, reducing the learning rate by half if the training loss does not improve after 200 epochs. We returned the mean estimate on the test data set over 10 runs. All experiments are run on NVIDIA RTX A6000 GPUs. The hyper-parameters are presented in Table 2 for MMG. Concerning the consistency tests, we independently train an autoencoder for each version of the mnist dataset with $r$ rows available.

### v2

@@ Abstract
Mutual information (MI) is one of the most general ways to measure relationships between random variables, but estimating this quantity for complex systems is challenging. Denoising diffusion models have recently set a new bar for density estimation, so it is natural to consider whether these methods could also be used to improve MI estimation. Using the recently introduced information-theoretic formulation of denoising diffusion models, we show the diffusion models can be used in a straightforward way to estimate MI. In particular, the MI corresponds to half the gap in the Minimum Mean Square Error (MMSE) between conditional and unconditional diffusion, integrated over all Signal-to-Noise-Ratios (SNRs) in the noising process. Our approach not only passes self-consistency tests but also outperforms traditional and score-based diffusion MI estimators. Furthermore, our method leverages adaptive importance sampling to achieve scalable MI estimation, while maintaining strong performance even when the MI is high. Code to reproduce experiments is at https://github.com/fengsxy/Diffusion_MI, and the unified MI library is at https://github.com/fengsxy/Diffusion-MI.
@@ 1 Introduction
Estimating Mutual Information (MI) from samples is a fundamental problem with widespread applications. Methods based on local density estimation (kraskov2004; pal2010estimation; gao2015efficient) have been displaced by variational approaches using neural networks (poole2019variational) to estimate lower bounds on MI (belghazi2018mutual; nguyen2010estimating). Unfortunately, these approaches may have sample complexity or variance which scale exponentially with the true MI (gao2015efficient; mcallester2020formal; oord2018representation). In practical scenarios, the MI between two variables is usually unknown, making reliable estimation challenging. To tackle this, song2019understanding introduced three self-consistency experiments designed to evaluate the robustness of MI estimators. For a more standardized and comprehensive evaluation, czyż2023normalevaluationmutualinformation developed a benchmark consisting of 40 synthetic datasets derived from diverse distributions, providing a consistent basis for assessing the performance of different MI estimators.
Denoising diffusion models jaschaneq; ddpm; vdm; itd have dramatically improved the modeling of complex distributions, igniting a new industry for AI art generation latent_diffusion; dalle2; imagen. Recently, minde introduced a MI estimator, MINDE, based on score-based diffusion models. This estimator not only achieved an 87.5% estimation success rate on the benchmark of czyż2023normalevaluationmutualinformation but also passed the self-consistency tests. However, its reliance on accurately approximating the log-density gradient can be a challenging intermediate step. This motivates a more direct formulation, connecting MI to the denoising objective itself rather than its gradient.
In this paper, we exploit the recently discovered connections between information theory and diffusion models to derive an elegant and effective approach to MI estimation (guo; itd). Our contributions are as follows:
  • We show that conditional density estimation and mutual information estimation can both be written exactly in terms of the global optimum of a denoising objective. Conditional density estimation corresponds to a gap between MMSEs (itd) for conditional and unconditional denoising diffusion models, and mutual information is the expected gap over all data, leading to the Mutual Information Estimation via the MMSE Gap in Diffusion (MMG) estimator.
@@ 2 Background
Let $p(\bm{z}_{\gamma}|{\bm{x}})$ be a Gaussian noise channel with $\bm{z}_{\gamma}=\sqrt{\gamma/(1+\gamma)}{\bm{x}}+\sqrt{1/(1+\gamma)}\bm{\epsilon}$ and $\bm{\epsilon}\sim\mathcal{N}(0,\mathbb{I})$, where $\gamma$ represents the Signal-to-Noise Ratio (SNR). The unknown data distribution is $p({\bm{x}})$. The MMSE refers to the minimum mean square error for recovering $\bm{x}$} in this noisy channel (itd).
@@
From itd we see that log likelihood can be written exactly in terms of an expression that depends only on the MMSE solution to the Gaussian denoising problem.
@@ 3.1 Mutual Information Estimation
Conditioning on $y$ can only decrease the MMSE wu2011functional, so the mutual information is non-negative as expected. While Eq. 7 appears to be novel, a version of Eq. 8 appeared in guo. This result holds for discrete or continuous data guo, or even mixed continuous and discrete, a particularly challenging problem gao2017estimating. We propose to use recent advances in denoising diffusion modeling to approximate the right-hand side to achieve better mutual information estimates.
@@ 3.2 Model Training
We parametrize $\hat{\bm{x}}_{\theta}(\bm{z}_{\gamma},\gamma,y)$ as some neural network that is trained with a mean square error loss to recover the signal, $\bm{x}$}, from noise and the auxiliary signal, $y$, across different SNRs, $\gamma$ (see Figure 2). In principle we should train a separate neural network that recovers data from noise without conditioning on $y$. However, we can borrow from existing architectures which train a single network for both conditional and unconditional denoising ho2021classifier. Conditional diffusion models have shown compelling results in glide; dalle2; imagen. During training we simply replace $y$ with a null value, $y=\emptyset$, some fraction of the time so that the same (conditional) network can learn to denoise in the absence of conditioning.
In practice, we can’t guarantee that the neural network finds the true global optimum, the MMSEs in Eq. 8. Because MMSEs appear with both signs in that expression, we can’t even guarantee that our network gives an upper or lower bound. While this is unfortunate, conventional wisdom suggests that sufficiently expressive neural networks do converge to global optima (du2019gradient; jacot2018neural).
@@ 4.1 Adaptive Importance Sampling
As in itd, we estimate integrals using importance weighted Monte Carlo estimators, with log-SNR values chosen from a logistic distribution with some location, $\mu$ and scale, $\sigma$. Since the shape of the integration area ($\Delta$ area in Fig. 2) is data-dependent, a fixed sampling distribution can be inefficient. For our adaptive variants, we therefore optimize this distribution for each specific task.
To find the data-adapted parameters $(\mu,\sigma)$, we employ a two-stage procedure. First, we train a preliminary model and analyze its conditional MMSE curve, $\text{MMSE}_{x|y}$. Inspired by the use of error landmarks in density estimation (itd), our heuristic identifies the critical transition region of this curve, where the denoiser becomes effective. Specifically, we define the parameters as follows (where $d$ is the data dimensionality):
@@ 4.2 The Orthogonal Principle
To further enhance estimator stability, we incorporate the orthogonality principle from kong2023interpretable. The principle states that for optimal denoisers, the gap in MMSE is precisely the expected squared distance between the conditional and unconditional estimators. This provides an alternative way to express the MI integrand. Formally, let $\hat{{\bm{x}}}(\bm{z}_{\gamma})=\mathbb{E}[{\bm{x}}|\bm{z}_{\gamma}]$ and $\hat{{\bm{x}}}(\bm{z}_{\gamma},y)=\mathbb{E}[{\bm{x}}|\bm{z}_{\gamma},y]$. The identity at a fixed SNR is:
@@
After expanding the left-hand side, the cross terms vanish due to the orthogonality principle kong2023interpretable, resulting in the expression on the right-hand side. This allows us to substitute the original integrand in our MI formula (Eq. 8) with the term on the right-hand side. The orthogonal MI estimator, in its practical importance-sampled form, is therefore:
@@ 5.1 MI Estimation Benchmark
We evaluate on the benchmark of (czyż2023normalevaluationmutualinformation), which combines base distributions (Uniform, Normal with dense or sparse correlation, and long-tailed Student-t) with MI-preserving nonlinear transformations (Half-Cube, Asinh, Swiss-roll, Spiral). This setup introduces high dimensionality, heavy tails, sparsity, and non-linear geometry. We compare MMG against neural estimators, such as mine (belghazi2018mine), Infonce (oord2019representation),nwj (nguyen2007neurips), doe (mcallester2020formal) and MINDE franzese2023minde.
@@
Table 1: Low MI estimation over 10 seeds using N = 10k test samples against ground truth (GT) (czyż2023normalevaluationmutualinformation). All methods were trained with 100k samples. Color indicates relative negative (red) and positive bias (blue). Blank entries indicate that an estimator experienced numerical instabilities. List of abbreviations ( Mn: Multinormal, St: Student-t, Nm: Normal, Hc: Half-cube, Sp: Spiral)
@@ 5.2 High MI Benchmark
To test the robustness and limits of our estimator, we extend the high-MI study from MINDE (minde), pushing their experimental setup from a range of MI $\leq$ 5 into the significantly more challenging regime of MI $\in[10,15]$. Following their protocol, we use a sparse $3\times 3$ Multinormal setup and its two MI-preserving transforms (Half-cube, Spiral). Figure 3 reports the mean $\pm$ one standard deviation over 10 seeds.
On the original and Half-cube cases (Fig. 3a&b), MMG-adaptive remains the most accurate as MI grows, while other estimators falter due to distinct sources of error. MINDE struggles significantly as its score-matching objective requires approximating the sharp, high-frequency score functions of high-MI distributions, a task where neural networks’ spectral bias leads to significant underestimation rahaman2019spectral. In contrast, our orthogonal variants exhibit a different limitation: a systematic conservative bias. This occurs because the orthogonal estimator measures the distance between neural network approximations of the denoisers, and the distance between these "smoothed-out" approximations is inherently smaller than the true distance between the optimal denoisers. This underestimation bias becomes more pronounced as the true MI gap grows larger.
@@ 5.3 Consistency Test
We conduct self-consistency tests inspired by song2019understanding to evaluate the properties of MMG using high-dimensional real-world data, specifically samples from the MNIST dataset (28×28 resolution) (mnist). Let $A$ represent an image and $B_{r}$ denote the image consisting of the top $r$ rows of $A$. These tests are designed to verify whether the estimators adhere to fundamental properties of MI through three subtasks: Baseline Test, Data-Processing Test, and Additivity Test for two independent images sampled from dataset. Ideally, as $r$ increases, $\frac{I(A;B_{r})}{I(A;B)}$ should monotonically approach 1, $\frac{I(A;[B_{r+k},B_{r}])}{I(A;B_{r+k})}$ should consistently equal 1, and $\frac{I([A^{1},A^{2}];[B^{1}_{r},B^{2}_{r}])}{I(A^{1};B^{1}_{r})}$ should consistently equal 2. To ensure a fair comparison with MINDE (minde), we aligned our experimental settings and parameters and tested MMG using five random seeds. The results of the three tests are presented in Figure 4. Overall, MMG performed well and successfully passed all tests.
@@ 6 Related Work
Estimating mutual information (MI) from samples is a central challenge in machine learning. Traditional non-parametric approaches, such as methods based on data binning or kernel density estimation (KDE), struggle with the curse of dimensionality. More advanced estimators based on k-nearest neighbors have shown significant improvements (kraskov2004; pal2010estimation), but can still be challenged by complex, high-dimensional data (gao2015efficient). Consequently, the modern paradigm is dominated by neural network-based approaches that optimize a variational bound on the MI.
@@
Most recent neural MI estimators are built upon variational lower bounds derived from f-divergences. The pioneering work, MINE (belghazi2018mutual), leverages the Donsker-Varadhan representation of the KL-divergence, spurring related estimators like the variance-reduced SMILE (song2019understanding) and DoE (mcallester2020formal). A particularly successful family is based on contrastive learning, where estimators like InfoNCE (oord2018representation) train a critic to distinguish between joint and marginal samples (poole2019variational). However, these methods face significant limitations: their sample complexity can scale exponentially with the true MI, often being capped by the logarithm of the batch size (mcallester2020formal). This difficulty can be viewed as a "density chasm": the marginal distribution $p(x)p(y)$ is often a poor proposal for the joint $p(x,y)$, leading to high-variance estimates, especially in high-MI settings (rhodes2020telescoping; brekelmans2021improving).
@@
Denoising diffusion models offer a natural and powerful framework to bridge this density chasm. They define a continuous process that transforms a complex data distribution into a simple tractable one, providing a path of intermediate distributions. The potential of this framework for MI estimation was recently demonstrated by MINDE (minde), which connects MI to the difference between conditional and unconditional score functions ($\nabla_{{\bm{x}}}\log p({\bm{x}})$). Our work, MMG, builds on a different and more direct connection (guo; itd). Instead of relying on learned score functions, we show that MI corresponds exactly to the integrated gap between the Minimum Mean Square Error (MMSE) of conditional and unconditional denoising. This formulation connects MI directly to the denoising objective itself, rather than its gradient, providing an elegant and potentially more robust pathway for estimation.
@@ Appendix A Derivation of the Orthogonal Principle
This section provides a brief derivation for the orthogonal principle (Equation 11) used in our estimator, following the original work of kong2023interpretable.
@@ Appendix C Implementation Details
We follow the implementation of franzese2023minde which uses stacked multi-layer perception (MLP) with skip connections. We adopt a simplified version of the same network architecture: this involves three Residual MLP blocks. We use the Adam optimizer (kingma2014adam) for training and Exponential moving average (EMA) with a momentum parameter $m=0.999$. We use the ReduceLROnPlateau scheduler with a patience of 200 epochs, reducing the learning rate by half if the training loss does not improve after 200 epochs. We returned the mean estimate on the test data set over 10 runs. All experiments are run on NVIDIA RTX A6000 GPUs. The hyper-parameters are presented in Table 2 for MMG. Concerning the consistency tests, we independently train an autoencoder for each version of the mnist dataset with $r$ rows available.